OpenStudy (anonymous):

y=-x+4 y=-6x+9 solution ?

OpenStudy (radar):

Since you were not told which method to use in solving "system" equations, why not use this logic. We know that y = y so lets put the two equations that equal y to equal each other like this: -x + 4 = -6x + 9 We can now solve for x -x + 6x+ 4= -6x + 6x + 9 by adding 6x to both sides getting 5x + 4 = 9 5x + 4 - 4= 9 - 4 by subtracting 4 from both sides getting 5x = 5 now divide both sides by 5 getting x = 1 We have now solved for x.

OpenStudy (radar):

To solve for y, simply substitute the value for x in either one of the 2 equations you were originally given. Any questions?

OpenStudy (radar):

Can you now solve for y? @raevenbennett